• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:54:59 on November 13, 1998

    Entry number 5310

    keyword=Beam Time Accounting
    Hall A

    11/12/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: R. Ransom
    Shift Workers: J. P. Chen, Z. Chai, G. Laveissiere

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 0 0 0
    Beam Available
    but Not in Use
    1 1 1 1 1 1 1 0 7
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 1 1
    Experiment
    Ready
    0 0 0 0 0 0 0 1 1
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    1 1 1 1 1 1 1 07
    Hall not
    ready reasons
    Target 1 1 1 1 1 1 1 0 7
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: Standby shift. 1600-2000 complete target work (installation of new cell) in the hall 2000-2100 laser door safer interlock/ work. 21:00-21:30 sweep, became conrolled access 21:30-23:00 target pumping up and study 23:00-2400 continue target pump up and started beam tuning

    Data Entry by: J. P. Chen